OpenStudy (anonymous):

The graph shows the system 3x + 6y = 18 and x + 2y = 6. What is the solution for the system? A. one solution B. no solution C. infinitely many solutions D. cannot be determined I say one solution

OpenStudy (anonymous):

infinitely many solutions

OpenStudy (anonymous):

Try (2, 2). That's a solution. But (4, 1) also works. Why? 3x + 6y = 18 x + 2y = 6 Multiply both sides of second equation by 3. 3x + 6y = 18 They're the same exact thing! If something works in the first one, it'll work for the second, so there are infinitely many solutions.